### v3

# RISCORE: Enhancing In-Context Riddle Solving in Language Models through Context-Reconstructed Example Augmentation

Abstract
Riddle-solving requires advanced reasoning skills, pushing LLMs to engage in abstract thinking and creative problem-solving, often revealing limitations in their cognitive abilities. In this paper, we examine the riddle-solving capabilities of LLMs using a multiple-choice format, exploring how different prompting techniques impact performance on riddles that demand diverse reasoning skills. To enhance results, we introduce RISCORE (RIddle Solving with COntext REcontruciton) a novel fully automated prompting method that generates and utilizes contextually reconstructed sentence-based puzzles in conjunction with the original examples to create few-shot exemplars. Our experiments demonstrate that RISCORE significantly improves the performance of language models in both vertical and lateral thinking tasks, surpassing traditional exemplar selection strategies across a variety of few-shot settings  [^1] Code for our experiments is available at: https://github.com/GiannisPana/RISCORE.
RISCORE: Enhancing In-Context Riddle Solving in Language Models through Context-Reconstructed Example Augmentation
Ioannis Panagiotopoulos, Giorgos Filandrianos, Maria Lymperaiou, Giorgos Stamou School of Electrical and Computer Engineering, AILS Laboratory National Technical University of Athens yiannispn@gmail.com, {geofila, marialymp}@islab.ntua.gr, gstam@cs.ntua.gr
1 Introduction
Reasoning in Natural Language Processing has been a field of increasing interest, especially since the surge of Large Language Models (LLMs) which showcase several reasoning shortcomings Yu et al. (2024); Qiao et al. (2023). The logical gap between humans and LLMs’ advanced reasoning is further exposed when probing the puzzle solving capabilities of such models Giadikiaroglou et al. (2024), denoting the necessity of intricate datasets Lin et al. (2021); Szomiu and Groza (2021); Zhang and Wan (2022) that test and enhance complex model reasoning. In terms of advanced reasoning skills, lateral thinking datasets such as BrainTeaser Jiang et al. (2023b) stress the creative reasoning capabilities of models, requiring them to defy obvious logical associations and approach a more abstract way of thinking.
At the same time, prompting is a widespread technique of extracting reasoning patterns from LLMs Qiao et al. (2023). Especially Chain-of-Thought (CoT) prompting Wei et al. (2023), as naturally designed to trigger the extraction of reasoning chains, is indeed proven to be one successful and viable prompting strategy in practice, even though its functionality is only applicable on larger models. Other promising techniques, such as in-context learning, are highly capable when the right data samples are utilized as exemplars, but fail to properly drive LLMs otherwise Dong et al. (2024). While similarity-driven exemplar selection serves as a well-established rule of thumb Liu et al. (2022); Qin et al. (2024); Dong et al. (2024), the notion of similarity becomes more obscure when concerning reasoning patterns: for example, the sentences A man shaves everyday, yet keeps his beard long and Tom attends class every day but doesn’t do any homework Jiang et al. (2023b) in fact represent the exact same reasoning pattern, despite being rather distant semantically.
Driven by this observation, we assume that by “stripping” a sentence from linguistic features, while focusing on pure reasoning patterns prevalent in sentences, we can provide the model with the critical information to unlock the desired underlying reasoning pattern. Sentences that represent the same reasoning pattern are provided as “context reconstruction” counterparts in BrainTeaser data Jiang et al. (2023b) or can be automatically generated using the provided method. Therefore, incorporating such reconstructed instances as few-shot exemplars, together with their original counterparts could serve as an alternative to existing few-shot strategies that respects reasoning similarity and analogy. An example of the method is provided in Figure 1. To this end, we contribute to the following:
  • We experimentally verify that providing a riddle along with its context reconstruction can enhance performance on both lateral and vertical thinking problems.
  • We propose RISCORE, a novel prompting method designed to enhance the in-context riddle-solving capabilities of LLMs. As a supplement to this, we introduce an algorithm for generating contextually reconstructed multiple-choice riddles.
  • We compare RISCORE against a wide array of popular prompting techniques, highlighting its effectiveness across several prompting alternatives and models of varying sizes.
Figure 1: Shows an example of few-shot (FS) prompting alongside the same method using a context-reconstructed example. The text in green represents the context reconstruction of Example 1, produced by the proposed method. The example in red was removed simply to maintain the same number of shots.
2 Related work
Reasoning with Language Models
Multiple reasoning senses have been studied over the years using language models, including commonsense Sap et al. (2020), arithmetic Luo et al. (2024), abductive Zhao et al. (2023), inductive Han et al. (2024), deductive Sanyal et al. (2022), analogical reasoning Sultan and Shahaf (2022) and others. Linear thinking processes that exploit rules and logic, termed as “vertical thinking”, have been previously explored within popular datasets Lin et al. (2021); Bisk et al. (2020), unveiling interesting reasoning patterns of language models. On the other hand, creative thinking has been widely underexplored and often purposefully excluded from reasoning benchmarks Speer et al. (2017); Sap et al. (2019), leading to a significant gap in literature, especially given the emergent capabilities that larger models present Wei et al. (2022). Puzzle solving closely lies to creative reasoning Giadikiaroglou et al. (2024); Lin et al. (2021); Zhang and Wan (2022), since out-of-the-box thinking is mainly required. Going one step further, overriding default presuppositions and associations occurring when reasoning leads to “lateral thinking” processes, leading to more tricky puzzles as demonstrated in the BrainTeaser dataset Jiang et al. (2023b) for the first time. In our work, we focus on probing vertical and lateral puzzle-solving reasoning abilities of language models.
LLMs and prompting
Discovering reasoning patterns in LLMs is often performed via various types of prompting Qiao et al. (2023). Zero-shot strategies harness simple but intuitive instructions, such as “Let’s think step-by-step” that successfully improve LLM reasoning Kojima et al. (2023); however, the search space of such magic prompts seems indefinite. Querying for intermediate reasoning steps was framed in the seminal Chain-of-Thought (CoT) prompting Wei et al. (2023), proving that it is sufficient to elicit reasoning in larger models. Nevertheless, the placement of demonstrations in few-shot prompting faces several challenges Dong et al. (2024); specifically, the way of selecting these exemplars themselves is prone to instabilities Lu et al. (2022); Min et al. (2022), proving the nonnegotiable need for developing optimal exemplar selection techniques. Retrieval based on similarity is established as a default selection technique Liu et al. (2022); Qin et al. (2024); Dong et al. (2024), with related enhancement optimizing ordering of selected exemplars Wu et al. (2023). With a focus on reasoning tasks, complexity of reasoning paths Fu et al. (2023) or diversity of exemplars Zhang et al. (2022) have already proven their merits, shifting the weight towards uncovering the patterns hidden in data themselves that are able to advance reasoning, rather than similarity-driven placement based on semantics. Inspired from the above, we craft exemplars that promote the hidden reasoning patterns rather than the semantics of related data samples, maintaining simple similarity-based retrieval for their placement; by outperforming other prompting techniques, we demonstrate that highlighting those reasoning patterns is adequate for advanced LLM reasoning, without further few-shot engineering.
3 Method
3.1 RISCORE prompting
Consider the following two riddles: R1: “A man shaves every day, yet keeps his beard long” and R2: “What has a beard but never needs to shave?”. While these riddles are semantically similar, follow the same structure, and refer to the same objects, their reasoning processes differ. In R1, where the answer is “A barber,” the word “beard” is used literally. In R2, where the answer is “A tree,” “beard” is used metaphorically, referring to the “beard” of certain trees, such as the “oak”. Thus, when using these riddles as context for a new riddle—such as “I plant seeds every day, yet don’t have a single plot”—the answer will depend on the interpretation of the phrase “plant seeds.” A metaphorical interpretation might lead to “A teacher,” while a literal one might suggest “A hydroponicist.” This illustrates how riddle context can significantly shape the reasoning process required for solving it.
In this example, however, instead of using R2, a contextually reconstructed version of R1R3: “Tom attends class every day but doesn’t do any homework”—would provide a clearer reasoning process to the model. Although R2 is semantically closer to R1, its metaphorical use of “beard” contrasts with the literal meaning in the original, potentially obscuring the reasoning pathway that the model should follow for solving a new riddle.
Building on the previous example, we propose the RISCORE prompting method, designed to enhance the in-context riddle-solving abilities of LLMs. The RISCORE method supplements each exemplar in FS learning with a contextually reconstructed version of itself. This preserves the original reasoning process while only altering its context. This approach enables the model to delineate a clear and coherent reasoning trajectory, which it can follow to effectively solve new riddles.
As illustrated in Figure 1, our approach builds on top of existing few-shot (FS) methods Dong et al. (2024); Wang et al. (2024); Sultan and Shahaf (2022), with RISCORE not involved in the FS selection process. The goal of our method is to augment FS samples with automatically generated context-reconstructed examples (detailed in Section 3.2). The addition of context reconstructed examples has, in most cases, proven to be more beneficial for the models’ performance, surpassing even the utility of real examples extracted from the dataset (see Section G).
3.2 Method for generating context reconstructions
In this section, we describe our approach for generating high-quality reconstructed Question-Answer pairs, which can be used as input to a model, alongside their counterparts, to enhance performance on lateral thinking problems. This process draws inspiration from the semi-automated pipeline introduced in the BrainTeaser paper Jiang et al. (2023b), but we extend it by fully leveraging the capabilities of LLMs to automate the procedure.
Step 1: Generation of Question-Answer pair
The first step involves generating one contextually reconstructed Question-Answer pair per selected instance, temporarily ignoring distractors. To achieve this, we provide the LLM with the riddle, the correct answer, and a system prompt that outlines the task. The prompt instructs the model to analyze the given Question-Answer pair, understand the riddle, identify the reasoning process that links the question to the answer, and then generate a similar riddle that follows the same reasoning process, along with a corresponding correct answer. Since the model is provided with both the riddle and the correct answer, understanding the riddle becomes easier than solving it independently Sultan and Shahaf (2022). To further refine this approach, we apply the method in both Zero-shot (ZS) and Few-shot (FS) settings, using preexisting pairs of original and contextually reconstructed questions from BrainTeaser. The prompts used in this process can be found in App. A.2. After generating the Question-Answer pair, a filtering procedure was conducted to ensure the quality of the instance and its alignment with the riddles in the dataset. This process involved applying rules concerning both the structure of the riddle and the answer. These rules are dataset-specific and can be adjusted to suit different datasets, ensuring adaptability across various contexts. Further analysis regarding this step can be found in App. A.1.
Step 2: Generation of the distractors
This process involves generating incorrect answers to complement the multiple-choice options. While this task may initially appear straightforward, it presents several significant challenges. First, the distractors must not only match the number of original options but also be guaranteed to be incorrect when compared to the correct answer. Additionally, they should not deviate excessively from the correct answer, as significant divergence could undermine the validity and challenge of the riddle. It is also essential to account for cases where the correct answer is “None of the above”, as observed in the BrainTeaser dataset. The length of the answers also plays a crucial role in this process. For instance, BrainTeaser includes mostly answers with more than four or five words, while RiddleSense provides typically single-word answers. These variations in length introduce additional challenges in crafting effective distractors. To address this, we propose two distinct methods for generating distractors: one for long and another for short distractors, respectively.
Generation of Long Distractors
In this approach, the original distractors from the riddle are inputted into a LLM alongside the new Question generated in Step 1. For each distractor (processed one at a time, except for “None of the above”), the model is prompted to modify its context by integrating it with the new riddle’s context. Regardless of the quality of this integration, the distractor is guaranteed to be incorrect, though lower-quality incorporation may slightly reduce the distractor’s overall effectiveness. If the original riddle’s answer differs from “None of the above” but the new riddle’s answer is, an additional distractor must be generated. In such cases, we simply prompt the LLM to generate a new distractor based on the generated Question-Answer pair. Details of the aforementioned step can be found in App. B.1.
Generation of Short Distractors
The second method for generating distractors is applied to datasets like RiddleSense, which contain single-word answers. In this scenario, modifying a single word like “water” to fit the riddle’s context presents a challenge. To address this, we first classify the generated Answer of Step 1 into a set of mutually exclusive categories (details in App. B.2), which informs the generation of distractors. For instance, in the riddle “I have a beard, but I never shave. What am I?” (where the correct answer is “A tree”), the system would classify the answer under the category “Nature”.
Next, to ensure the distractor is incorrect, we break the riddle into smaller phrases. We also detect the presence of interrogative words, and if the question is purely descriptive without a direct query, we append the phrase “What am I?” at the end. For the previous example, we split the riddle into two parts: “I have a beard, what am I?” and “I never shave. What am I?” These phrases are then provided as inputs to the LLM, and only answers that differ from the correct one are retained. Since the LLM has access only to a phrase, we expect the generated answer to be incorrect. However, this is not always the case. In the example above, the response to “I have a beard, what am I?” might be “An oak,” which is also a valid answer for the riddle.
To address this issue, we leverage the classification of the Answer to guide distractor generation. Specifically, we provide the LLM with two closely related categories, excluding the correct answer’s category (e.g., “person” and “place”). The model is then prompted, through a system-user instruction, to generate a valid answer within the given context and category. This approach ensures that the generated distractors are both incorrect and contextually relevant, thereby preserving the difficulty of the riddle. After applying a filtering stage to ensure the distractors are distinct, we use WordNet Miller (1992) to augment them if their quantity is insufficient. Further details can be found in App. B.2. Additionally, for both methods, the same filtering procedure applied to the correct answer is also used to ensure the quality of the distractors.
Step 3: Creation of the Riddle
In the final step of this pipeline, we use the Question-Answer pair, randomly shuffle the distractor set, select the required number of distractors, and place the correct answer in a randomly assigned position. More details regarding the Prompting process for our method can be found in App. C.
4 Experiments
We evaluate various models using different datasets with our method and compare their performance against multiple baselines.
4.1 Datasets
For our experiments, we utilized two different datasets, namely BrainTeaser and RiddleSense. A comprehensive analysis of the datasets used in our experiments can be found in App D.
We selected BrainTeaser primarily for two key reasons. First, BrainTeaser provides high-quality, manually crafted context reconstructions, which establish an upper bound on the model’s performance when these are incorporated into the input. This upper bound also serves as a benchmark for evaluating the quality of riddles generated by the automated method. Second, the dataset consists of riddles that test the lateral thinking abilities of the models, a reasoning process that has proven as a challenging cognitive process for LLMs Giadikiaroglou et al. (2024).
Additionally, experiments were conducted using the RiddleSense dataset, which emphasizes vertical thinking, to evaluate the method’s applicability across different types of riddles. Given that hand-crafted context reconstructions are not available for RiddleSense, the automated method was utilized exclusively in this set of experiments.
4.2 Baselines
In our study, we used various prompting techniques as baselines, including zero-shot (ZS), few-shot (FS), and Chain of Thought (CoT) methods. A detailed analysis of each technique can be found in the App. E.
Chain-of-thought (CoT_ZS)
In this experiment, we prompt the model to answer the multiple-choice riddle in a step-by-step manner, following the method proposed in Wei et al. (2023), without including any examples. This approach allows us to assess the models’ riddle-solving abilities in a zero-shot setting.
Few-Shot (FS)
To assess the influence of example-based learning, we incorporated few-shot prompting with varying numbers of examples: 2, 4, and 8 shots. We limited our tests to 8 shots since it has been proven that reasoning abilities of LLMs are deteriorating with longer inputs Levy et al. (2024). Two distinct strategies were used for selecting these examples: random selection, referred to as Rand, and selecting riddles with minimal semantic distance from the test riddle using a language model Zhang et al. (2024) [^2] https://huggingface.co/Alibaba-NLP/gte-large-en-v1.5, referred to as Sim.
Few-shot with CoT Explanations (CoT_FS)
A straightforward way to aid the model in understanding an exemplar is by providing a supplementary explanation. In this experiment, we explored this approach by offering explanations alongside the few-shot exemplars. This allowed us to evaluate whether these crafted explanations could improve the models’ ability to comprehend and follow the required reasoning process. The explanations were generated using a semi-automated process. First, for each exemplar in the FS setting, we used ChatGPT [^3] Specifically, we used version gpt-3.5-turbo-0125. to generate an explanation for the riddle based on the correct answer, following the method proposed in Wei et al. (2023). These explanations were then manually curated to ensure alignment with human interpretations of the reasoning process required for the riddle. Due to the labor-intensive process of annotating the explanations, we conducted this experiment using only a randomly selected sampling strategy. Additionally, the results indicate that the explanations do not significantly aid the models in solving riddles (see Section 5 for further details).
4.3 RISCORE
RISCORE is applied as follows. For each example used in the few-shot (FS) setting, the corresponding contextually reconstructed riddle is appended to the input. For a fair comparison, the number of shots refers to the total number of examples included in the prompt, regardless of whether they are new riddles or context reconstructed examples. Thus, for example RISCORE with 4 shots has the same 2 examples used in 2-shots FS Sim technique augmented with the context reconstruction of them, meaning that keeps only half of the examples drawn from the original dataset and the remainder consisting of reconstructed examples. The experiments were conducted with 2, 4, and 8-shot settings, meaning that 1, 2, and 4 exemplars, respectively, were drawn from the dataset, with the rest being their context reconstructions.
We primarily utilized the Llama3-8B and Llama3-70B model to generate both Question-Answer pairs and distractors. The Question-Answer pairs were produced in both zero-shot and few-shot settings.
Lastly, in our experiments, we differentiate between RISCOREm, which utilizes manually created reconstructions (when available), and RISCORE, which employs the fully automated method for generating reconstructions. The prompt structure for the system-user interaction in this method is the same as in the FS method outlined in App. E. The sole distinction is in the examples provided.
Method N. Llama3-70B Mistral-8x7B Llama3-8B Mistral-7B Qwen2-7B
CoT_ZS 0 0.725 0.550 0.633 0.450 0.458
Randomly Selected Shots
2 0.758 0.617 0.633 0.475 0.608
CoT_FS 4 0.683 0.583 0.608 0.508 0.650
8 0.708 0.642 0.658 0.508 0.667
2 0.775 0.617 0.633 0.517 0.642
FS 4 0.808 0.683 0.642 0.483 0.608
8 0.775 0.617 0.675 0.483 0.642
2 0.783 0.625 0.667 0.458 0.608
RISCOREm 4 0.758 0.617 0.675 0.517 0.625
8 0.800 0.650 0.667 0.400 0.592
Semantically Similar Shots
2 0.825 0.692 0.700 0.517 0.600
FS 4 0.792 0.683 0.717 0.458 0.633
8 0.783 0.667 0.767 0.533 0.650
2 0.783 0.675 0.767 0.483 0.667
RISCOREm 4 0.833 0.708 0.742 0.567 0.642
8 0.808 0.708 0.758 0.550 0.667
Table 1: Model’s performance for Brain Teaser using baselines and RISCOREm prompting. The best overall results are highlighted in bold. More detailed results can be found in Tables 6 and 7.
Method N. Llama3-70B Mistral-8x7B Llama3-8B Mistral-7B Qwen2-7B
Llama3-70B zeroshot for QA & Llama3-8B distractors
2 0.792 0.667 0.625 0.492 0.625
RISCORE 4 0.792 0.642 0.675 0.467 0.625
8 0.808 0.683 0.700 0.475 0.642
Llama3-70B fewshot for QA & Llama3-8B distractors
2 0.750 0.675 0.683 0.475 0.625
RISCORE 4 0.792 0.650 0.658 0.558 0.658
8 0.808 0.675 0.742 0.517 0.658
Llama3-70B fewshot for QA & Llama3-70B distractors
2 0.783 0.667 0.683 0.500 0.617
RISCORE 4 0.792 0.642 0.667 0.508 0.617
8 0.767 0.683 0.700 0.500 0.617
Table 2: Model’s performance for Brain Teaser using RISCORE prompting. Similarity-based selection was employed for choosing all the exemplars. Results that surpass the FS method with semantically similar examples, using the same number of shots, are underlined. More detailed results can be found in Tables 7 and 8.
4.4 Models
To this end, we test various models of different scales. Specifically, we examine Llama3 with 8 and 70 billion parameters respectively AI@Meta (2024), Mistral-7B and Mistral-8x7B Jiang et al. (2023a), as well as Qwen2-7B Yang et al. (2024). This diverse set of models allows us to investigate the impact of contextually reconstructed examples on models with varying sizes and parameterizations. Further details regarding the model selection process and technical specifications can be found in App. F.
5 Results
In this section, we present and analyze the results of our experiments. Additional results, utilizing different metrics and hyperparameter settings, are available in App. G.
5.1 BrainTeaser
Table 2 presents the performance outcomes of various prompting techniques applied to the BrainTeaser dataset. Notably, the CoT_FS method underperforms compared to few-shot techniques, even when supplied with examples that include manually crafted explanations. This trend holds regardless of the model size utilized. Conversely, there is a significant improvement when the few-shot approach incorporates the provided “context reconstructed” riddles alongside the original riddles. This suggests that context reconstructed riddles effectively guide the model in tracing the appropriate reasoning path. Remarkably, the benefit of using these reconstructed contexts is more pronounced in smaller models compared to larger ones of the same architecture. For instance, the improvement in the Llama3-8B model is substantial: while the 4-shot technique using semantic similarity achieves a score of 0.717, augmenting this approach with four context reconstructed examples (totaling eight shots) enhances performance to 0.758—an improvement of approximately 5.7%. For the larger model, this increase is from 0.792 to 0.808, leading to an improvement of approximately 2.02%. The same pattern is more evident to the Mistral-7B, when for 4-shot we notice an increase of 9.2% while for Mixtral-8x7B we have an increase of 2.5%. A similar trend can be observed in Qwen2-7B, where the 4-shot setting with semantically similar examples performs approximately 3.4% lower than the augmented 8-shot exemplars, which include the same examples along with their context reconstructions, resulting in a total of 8 examples.
At this point, it is worth noting that the performance improvement does not originate solely from the increased number of shots but also from the quality of the shots added. This is evident from the fact that the best results on the dataset are achieved when we include examples with their context reconstruction riddles. Furthermore, by comparing performance while keeping the number of shots constant, we observe that cases using the reconstructed riddles perform better, especially when the number of shots is more than two. In the latter case, the addition of different examples seems more helpful to the model in properly understanding the task, since including only a single example with its contextual reconstruct makes the prompt overly specific.
Using the automated method
Table 2 presents the performance of RISCORE using the automated method for generating context reconstructed examples. As expected, the using the automated method yields a smaller improvement in model performance compared to the manually curated examples provided in the BrainTeaser dataset. However, the performance of all models was consistently improved by the addition of context reconstructed riddles, even when using the fully automated algorithm for their generation. For example, for the Llama3-8B model, the performance using 4 shots from the dataset is 0.717, and by using the same shots along with their reconstructions generated by the automated method (8 shots in total), the performance increases to 0.742. The same trend is evident across all relatively smaller models, where augmenting the 4-shot examples selected based on semantic similarity with their automated context pairs yields notable performance improvements. Specifically, in the case of Qwen2-7B, the 4-shot setting augmented with corresponding context exemplars results in a 2.5% performance improvement. In Mistral-7B, the improvement is even more pronounced, with performance increasing by 10% from a baseline of 0.458 in the best-case scenario. Importantly, Table 2 shows that in most cases (underlined), it is even better to use the automated method to generate additional examples compared to adding the same number of examples from the dataset. For instance, the performance of Llama3-70B with 8 shots from the dataset is 0.783, whereas using only 4 examples from the dataset and automatically creating the remaining 4 using the context reconstruction method results in a performance of 0.808.
Method N. Llama3-70B Mistral-8x7B Llama3-8B Mistral-7B Qwen2-7B
CoT_ZS 0 0.775 0.675 0.619 0.589 0.608
Randomly Selected Shots
2 0.789 0.692 0.625 0.594 0.667
CoT_FS 4 0.783 0.686 0.672 0.603 0.656
8 0.783 0.697 0.658 0.597 0.625
2 0.769 0.706 0.672 0.586 0.689
FS 4 0.772 0.719 0.639 0.586 0.683
8 0.800 0.711 0.672 0.586 0.700
Semantically Similar Shots
2 0.792 0.714 0.706 0.608 0.722
FS 4 0.817 0.692 0.711 0.633 0.714
8 0.800 0.675 0.681 0.611 0.731
Table 3: Model’s Performance for RiddleSense using baseline techniques. The best results overall are in bold. More detailed results can be found in Table 9.
Method
N. Llama3-70B Mistral-8x7B Llama3-8B Mistral-7B Qwen2-7B
Llama3-70B fewshot for QA & Llama3-70B distractors
2 0.792 0.672 0.692 0.600 0.697
RISCORE 4 0.783 0.689 0.722 0.600 0.717
8 0.789 0.700 0.708 0.597 0.731
Llama3-70B fewshot for QA & Llama3-8B distractors
2 0.786 0.719 0.681 0.603 0.681
RISCORE 4 0.789 0.686 0.686 0.606 0.697
8 0.775 0.689 0.706 0.617 0.719
Llama3-8B zeroshot for QA & Llama3-8B distractors
2 0.792 0.681 0.689 0.589 0.694
RISCORE 4 0.778 0.714 0.700 0.600 0.683
8 0.806 0.689 0.686 0.614 0.689
Table 4: Model’s performance for RiddleSense using RISCORE prompting. Similarity-based selection was employed for choosing all the exemplars. Results that surpass the FS method with semantically similar examples, using the same number of shots, are underlined. More detailed results can be found in Table 10.
5.2 RiddleSense
In this dataset, RISCORE can only be applied to automatically generated examples because the dataset’s format lacks context reconstructions for its questions.
Table 4 presents the results of the baseline techniques using various models in the RiddleSense dataset. Once again, the results confirm that the few-shot technique, utilizing contextually similar exemplars for in-context learning, consistently delivers the best performance across all tested models.
In Table 4, we present the RiddleSense results using the proposed method for context reconstruction of each input. A clear trend emerges when comparing a simple 8-shot exemplar selection based on semantic similarity with the 8-shot setting of our method, where the top 4 semantically similar examples are augmented with our generated contextual pairs. Notably, the results show that our method consistently outperforms the standard 8-shot exemplar approach, demonstrating a significant improvement in model performance across various instances. An example of this trend is observed with the Llama3-8B model, where our method scores 0.708, approximately 2% higher than the few-shot setting based on semantic similarity, which achieves a score of 0.681. The same pattern is evident when comparing the two 4-shot settings. We achieve similar or marginally better accuracy using a total of just 4 examples—two original and two generated contextual reconstructions. This underlines the effectiveness of integrating contextually reconstructed pairs in enhancing model accuracy. We do not achieve a significant performance boost; however, we attain similar or marginally better results while relying on less grounded knowledge. This demonstrates the efficiency of our method, as it maintains comparable performance with fewer, yet strategically selected, exemplars.
5.3 Quality of context reconstructed riddles
To generate context reconstructed riddles, we used Llama3 models with 8 billion and 70 billion parameters in both few-shot (FS) and zero-shot (ZS) settings. We found that the Llama3-8B model struggled to produce high-quality Question-Answer pairs for the BrainTeaser dataset and was therefore not used in RISCORE. This difficulty likely arises from the BrainTeaser dataset’s demand for lateral thinking, which is particularly challenging for the smaller model. The Question-Answer pairs are essential, and if their quality is insufficient—as observed with Llama3-8B on the BrainTeaser dataset—high-quality distractors alone cannot compensate for this deficiency. However, for vertical thinking riddles, the smaller model effectively generates riddles even in the ZS setting, which, when used in the few-shot (FS) setting, can lead to increased performance compared to real examples drawn from the dataset.
6 Conclusion
We explore the riddle-solving capabilities of LLMs through multiple-choice formats, investigating how different prompts affect riddle performance requiring diverse reasoning. Our findings suggest that in few-shot settings, augmenting exemplars with contextually reconstructed examples improves performance. We introduce a novel prompting method, RISCORE, validated using the BrainTeaser dataset featuring manually crafted contextual riddles. Inspired by these results, we provide a method for automatically generating contextual reconstructions for multiple-choice riddles, demonstrating that RISCORE enhances LLMs’ abilities in lateral and vertical thinking tasks.
Limitations
This method has some limitations. For instance, the placement of exemplars is based on semantic similarity, although cues unrelated semantically might be more crucial for reasoning, such as reasoning path similarity or diversity Zhang et al. (2022). To address this, future work will employ more advanced methods for selecting exemplars for few-shot settings, like those suggested Sultan and Shahaf (2022). Additionally, the current study, having tested only two datasets, limits the broader impact of the conclusions since outcomes might vary with other datasets. Future steps will include testing the RISCORE prompting method with additional datasets across various tasks Zhang and Wan (2022). Finally, our experimentation is exclusively focused on the English language, therefore we cannot conclude whether our findings are applicable on other languages.
Ethical Considerations
There are no considerable risks associated with our work.
Appendix A Question-Answer Pairs Generation
A.1 Process details
This section of our pipeline involves several steps, and subsequent steps after the initial one vary based on the dataset format.
Initially, we need to create contextual reconstructed Question-Answer pairs, temporarily disregarding distractors for both datasets. To generate these Question-Answer pairs, we can leverage the capabilities of large language models (LLMs) in two ways. First, we use a system-user prompt template to provide specific instructions to the model, guiding it through the process of producing a new Question-Answer pair. This approach maintains the underlying commonsense premise of the original question while altering both the question and the answer to fit a new situational context. This method operates in a zero-shot setting, meaning the model generates the new pairs without prior examples or training specific to the task. The second approach is quite similar but leverages the pre-existing pairs of original and contextually reconstructed questions from BrainTeaser. In this few-shot setting, we use the same instructions as in the zero-shot approach, but we enhance the model’s understanding by providing in-context examples. Specifically, we include a pair consisting of an original Question-Answer and its corresponding contextually reconstructed version. This helps the model better grasp the task at hand. It is important to note that these few-shot examples are of high quality, given their manual construction and careful curation. In the few-shot approach, the provided examples are static. It is important to highlight that, in BrainTeaser, when selecting examples for few-shot exemplars from the training dataset, we ensure that the examples chosen are not selected for reconstruction. This avoids any overlap between the few-shot exemplars and the examples for which we are requesting a context change, ensuring that the given pair examples are distinct from the requested ones. The corresponding system and user prompts regarding the above tasks can be found in App. A.2.
Models utilized
In this approach, we experiment with two models to assess the impact of model size on the quality of contextual reconstructions and the overall performance across both datasets. We use a relatively small model, Llama3-8B AI@Meta (2024), and a larger model from the same family, Llama3-70B AI@Meta (2024). These models were selected due to their strong performance across a diverse range of tasks, outperforming other models of comparable size. By incorporating both a smaller model, Llama3-8B, and a larger model, Llama3-70B, we aim to systematically investigate the effect of model size on the quality of contextual reconstructions and its overall impact on performance. This exploration allows us to gain deeper insights into how model capacity influences the effectiveness of the generated examples across different datasets.
Quality Assurance for Generated Text
After generating the contextual reconstructed Question-Answer pairs for all initially selected examples from the training set, we perform a small but necessary quality control to ensure that the generated pairs adhere to several high-level criteria. For the BrainTeaser dataset, the filtering process is more flexible, as the questions and answers can vary in length. The only constraint is that the questions must have a minimum length of 7 words, while the answers can be as long as necessary. As a result, the number of generated pairs filtered out is relatively low, around 2%, ensuring that the majority of reconstructions meet the quality criteria. In contrast, the RiddleSense dataset requires stricter limitations due to its typically shorter answers. For this dataset, questions must be at least 6 words in length, while answers must not exceed 7 words, reflecting the dataset’s concise nature. As a result of these stricter rules, approximately 10% of the generated reconstructions for the RiddleSense dataset are filtered out for not meeting the required standards. It is important to note that the minimum question length requirement is implemented to ensure that the generated questions maintain a sufficient level of complexity and depth. Questions with fewer than 6 or 7 words are unlikely to provide the necessary detail to establish a well-structured and coherent riddle. This limitation helps preserve the quality and integrity of the riddles by ensuring that they are adequately framed to challenge and engage the model effectively.
A.2 Prompts
First, we will present the zero-shot prompt configuration used for generating the contextually reconstructed Question-Answer pairs.
  System Prompt:
You are an expert in context reconstruction. Your task is to receive a question along with its correct answer and adapt them to a new scenario while maintaining the misleading commonsense premise.
Please follow these steps:
- First, you will receive an unsolved riddle along with five answer options. Analyze the given setting and identify the connection between the question and its correct answer.
- Modify the original question and correct answer to fit a different situational context, ensuring that the underlying logic and relationship between them are preserved.
- Ensure that both the new question and the new correct answer are distinct from the originals.
User Prompt:
Question: ‘‘‘{QUESTION}‘‘‘
Correct answer: ‘‘‘{ANSWER}‘‘‘
  Next, with minor modifications, we introduce the few-shot prompt configuration that follows the same logic but incorporates in-context examples to guide the model’s responses.
  System Prompt:
You are an expert in context reconstruction. Your task is to receive a question along with its correct answer and adapt them to a new scenario while maintaining the misleading commonsense premise.
Please follow these steps:
First, review an example provided with its context reconstruction, which illustrates the type of transformation you will need to perform.
Next, you will receive an unsolved riddle along with five answer options. Analyze the given setting and identify the connection between the question and its correct answer.
Modify the original question and correct answer to fit a different situational context, ensuring that the underlying logic and relationship between them are preserved.
Ensure that both the new question and the new correct answer are distinct from the originals.
User Prompt:
{EXAMPLES}
Adapt the following riddle - answer pair while taking into consideration the examples above regarding context reconstruction:
Question: ‘‘‘{QUESTION}‘‘‘
Correct answer: ‘‘‘{ANSWER}‘‘‘
  In the above setting, it is understood that the values of the EXAMPLES represent pairs of answered Question-Answer examples: the original and its contextually reconstructed counterpart, both sourced from the BrainTeaser dataset.
Appendix B Distractor Creation
B.1 BrainTeaser
In this process, we leverage the capabilities of large language models (LLMs) to understand and rephrase context in order to generate distractors. We use two different pipelines to produce at least three distractors, ensuring coverage of various potential situations. 1The first approach involves prompting the model with a system-user prompt that instructs it to analyze the given Question-Answer pair. The model is tasked with understanding the riddle, identifying the reasoning process that links the question to the answer, and then suggesting a distractor based on the more challenging or deceptive aspects of the concept. This approach yields one of the three required distractors. 2In the second approach, the process is more intricate and can result in distractors of questionable quality. We prompt the model with a system-user prompt, providing it with the reconstructed question without its answer, along with the original question’s incorrect distractors, after removing the option “None of the above”. The model is then tasked with modifying the concept of the given distractors by incorporating elements from the setting described in the question. Importantly, the correct answer is not provided to ensure that the generated distractors are distinctly different from the correct answer and effectively capture a varied interpretation of the question’s context. Certainly, there are instances where the model generates sub-optimal contexts for the distractors provided. Despite these cases, the distractors remain incorrect and serve their purpose, though they may not always be sufficiently challenging. Our experiments indicate that when the model is given both the original and the contextual reconstructed examples, the minor issue of lower-quality distractors does not significantly impact overall performance. Now, we have also created two new distractors that are somewhat contextually relevant to the setting, adding an additional layer of coherence to the generated options. This ensures that the distractors are not only incorrect but also related to the underlying premise of the riddle, enhancing the overall quality of the multiple-choice options. To create the final dataset, we randomly select two of the three generated distractors and shuffle them with the correct answer in random order. Finally, we append the option “None of the above” as the last choice. In cases where “None of the above” is the correct answer, we use all three generated distractors. With this approach, our dataset is prepared and ready for use.
Prompts used
First, we will provide the system-user prompt used to task the model with understanding the riddle, identifying the reasoning process linking the question to the answer, and then suggesting a distractor based on the more challenging or deceptive aspects of the concept.
  System Prompt:
Your task is to act as a concept grasper. You will be given a riddle and its correct answer.
Your goal is to understand the connection between the riddle and the correct answer, focusing on the tricky parts. Based on these tricky aspects, propose a plausible wrong answer that someone might give.
The wrong answer should be short, concise, and limited to one sentence.
- Riddle:
- Correct Answer:
Response format:
- Wrong Answer:
User Prompt:
- Riddle: {QUESTION}
- Correct Answer: {ANSWER}
  Now we will provide the system-user prompt used for the second method.
  System Prompt:
You will be given a sentence without context and then provided with a specific context.
Your task is to rewrite the sentence so that it aligns with the given context, while keeping it as close as possible to the original meaning.
The purpose is to adapt the sentence to the context, not to answer any questions related to the context.
- Sentence (out of context):
- Context:
Response format:
- Sentence:
User Prompt:
- Sentence (out of context): {ORI_CHOICE}
- Context:{QUESTION}
  Here, the value of ORI_CHOICE refers to the distractors from the original instance, excluding the "None of the above" option. We create prompts for the model with each of these distractors individually.
B.2 Riddlsense
In this setting, the approach is fundamentally different. The answers and distractors in the original dataset are primarily one-word responses, while the questions feature detailed settings with punctuation, conjunctions, and more complex structures. To handle this, we also generate distractors using two distinct pipelines tailored to this format.
First pipeline
The first pipeline involves a more granular approach by splitting the contextual reconstructed question into subphrases based on punctuation or conjunctions. If this yields fewer than three distinct subphrases, we further split the sentence at the position of the word “and” to create additional segments. We also detect the presence of interrogative words, and if the question is purely descriptive without any direct question, we append the phrase “What am I?” at the end. This is not chosen arbitrarily but follows the standard structure of many riddles in this dataset, where "What am I?" is the common question leading to single-word answers. Now that we have broken the riddle into sub-phrases and appended the appropriate question, we can prompt the model to generate a wrong answer for each sub-phrase concatenated with the question. This method ensures that the distractors align with different parts of the riddle’s setting. However, since some sub-phrases may contain key ideas central to the riddle, the model may still generate answers too similar to the correct one. To address this issue, we incorporate an additional intermediate step utilizing zero-shot classification. Specifically, we use the facebook/bart-large-mnli model 444facebook/bart-large-mnli Lewis et al. (2019) provided by Hugging Face. This model classifies the correct answer into one of eight general categories: ’food’, ’person’, ’object’, ’animal’, ’nature’, ’time’, ’place’, ’concept’. These categories are chosen to be mutually exclusive to avoid overlap. In our approach, we first predict the category of the correct answer using this model. We then leverage this classification to guide the generation of distractors. For each sub-phrase concatenated with the question, we provide the LLM with the two most similar categories (excluding the correct answer’s category). The model is prompted with a system-user instruction to produce a correct answer of the given setting in the given category. This method ensures that the generated distractors are not only incorrect but also contextually aligned with the riddle, thus maintaining the challenge for the model. After applying filtering to ensure the distractors are distinct and relevant, we have produced several distractors. The LLMs utilized in this pipeline include the previously mentioned Llama3-8B AI@Meta (2024) and its larger counterpart, Llama3-70B AI@Meta (2024).
Second pipeline
In cases where the distractors produced by the above pipeline are insufficient, we use WordNet Cognitive Science Laboratory, Princeton University (2006) to augment our distractor set. For each generated distractor, or if necessary, for the distractors from the original question, we retrieve synonyms and hyponyms from WordNet. These additional terms are then included as potential distractors. After compiling these distractors, we randomly select four of them and add the correct answer in a random order to complete the set. This approach ensures that we have a diverse and comprehensive set of distractors for each question. To ensure the quality of our distractors, we impose a limitation that at least two of the required four distractors must be generated using the first approach. This is due to the fact that the distractors generated through WordNet augmentation tend to be of inferior quality compared to those produced directly by the model. If this requirement is not met, we skip the particular train set instance for producing a contextual reconstruction. This approach helps maintain the overall quality and relevance of the distractors in our dataset.
Prompts used
We will present the system-user prompt configuration used in the first pipeline. In this approach, for each sub-phrase concatenated with the question, we provide the model with two categories, excluding the correct answer’s category. For each category, a separate system-user prompt is issued to instruct the model to generate a correct answer that fits the given setting within the specified category.
  System Prompt:
Task: Provide a concise, relevant answer to the given question within the specified category.
Constraints:
- The answer should not exceed three words.
- Follow the exact format provided below.
Response Format:
Answer: …
User Prompt:
Question:  ‘‘‘{QUESTION}‘‘‘
Category: {CATEGORY}
Appendix C Pairs Prompting
We have successfully created the desired contextual reconstructions for both datasets. However, an issue remains: some of the originally selected examples do not have corresponding reconstructed examples due to the quality control filtering process.
To address this issue, we first use the original most semantically similar examples as in-context learning exemplars, appending their corresponding automatically generated contextual pairs. If these exemplars are not sufficient for the required settings of two, four, or eight (i.e., RISCORE), we employ a more structured approach rather than adding examples randomly. 1 We begin by generating embeddings for the set of original examples and their contextual reconstructions that have not already been included in the current exemplars. Using cosine similarity, we then identify the most similar examples from this set. Importantly, these similar examples might not be part of the original training set but could be among the reconstructed examples. 2 We select the most similar examples and pair each with its corresponding pair, ensuring that we add either the original or the reconstructed example as needed. This process is repeated until the number of exemplars meets the required quantity.
Appendix D Dataset Description
In this work we are using two different datasets of multiple QA format. 1 The first one is the BrainTeaser dataset, which consists of two sub-tasks regarding both lateral thinking challenges. In addition to the original puzzles, the dataset includes adversarial subsets created by manually modifying the original BrainTeaser while preserving their reasoning paths. The original data were perturbed in two ways: First, there is a semantic reconstruction of each original question, ensuring that the answers and distractors remain unchanged. Second, the original data undergoes context reconstruction, where the reasoning path is preserved, but the brain teaser is rephrased to describe a new situational context. Each question offers four possible choices, with the last option always being “None of the above”. It is important to note that our experiments are focused exclusively on the sentence puzzle (SP) setting. 2 The second dataset used in our experiments is RiddleSense. This dataset features riddle-style questions that require sophisticated commonsense reasoning and a strong grasp of figurative language to answer accurately. It is structured as a multiple-choice question answering task, with riddles that test the model’s capacity to navigate and interpret nuanced commonsense scenarios. The reasoning challenges posed by RiddleSense are closely aligned with those found in BrainTeaser. RiddleSense offers one correct answer and four distractors for each riddle. This contrasts with BrainTeaser, which includes a “None of the above” option as the final choice, thereby introducing a more challenging setting.
D.1 Data statistics
Format of Datasets
The BrainTeaser dataset is divided into three splits: train, development, and a hidden test set used for evaluation. Although RiddleSense follows a similar format, we did not have direct access to its hidden test set. As a result, we utilized the development data split of RiddleSense for our experiments.
Prepossessing and filtering
Due to data leakage across the two datasets, we were required to exclude semantically similar examples from the test split utilized. This leakage is understandable, as the sources for English riddles are limited, and the number of unique riddles is finite. After identifying semantically similar riddles, we retained only one instance within the datasets, prioritizing BrainTeaser due to the higher quality of its questions and distractors, which were manually crafted. Additionally, hardware limitations made processing the full RiddleSense unfeasible. To manage this, we split its test set into two halves. To ensure the fairness and accuracy of this division, we evaluated various techniques on both halves. Following the completion of these procedures, we present the dataset statistics in the table below. Statistics regarding the datasets are provided in Table  5. More information regarding this process can be found in D.3.
Dataset Train Dev Test
BrainTeaser - SP 507(169x3) 120(40x3) 120(40x3)
RiddleSense (initial) 1021
RiddleSense (filtered) 3510 720
RiddleSense (sampled 50%) 360
Table 5: Data statistics
D.2 Evaluation metrics
Given that our task is formatted as a multiple-choice question-answering problem, the evaluation metric employed will be accuracy. This metric will provide a measure of the proportion of correctly answered questions out of the total number of questions presented. Although accuracy is a straightforward metric, it is effective in providing useful insights and serves as a reliable measure for evaluation in this context.
1
For the BrainTeaser task, we not only track the overall accuracy but also monitor the accuracy for each specific type of instance, including original, semantic, and context reconstructions. This detailed tracking is feasible because the data in each set are balanced, with each original instance having corresponding semantic and context variations. This approach allows us to evaluate the model’s proficiency in handling the same reasoning path under different conditions. The total accuracy is then calculated as the mean of these three individual metrics. We also track group-based accuracy, where a "group" refers to either two questions (original and semantically reconstructed) or three questions (original, semantically reconstructed and contextually reconstructed). This metric assesses the model’s performance when all instances within a group are answered correctly. Group-based accuracy provides a broader perspective on the model’s ability to handle these lateral thinking challenges across different types of question reconstructions. This format of evaluation metrics was both provided and requested by the creators of the dataset in an open Kaggle competition, and we are adopting it for our analysis.
2
Unfortunately, the simple data row format in RiddleSense restricts our ability to track more detailed metrics comprehensively, limiting us to evaluating only the overall accuracy for this dataset.
D.3 Dataset preprossessing
Our goal in this process is to ensure that no question from the RiddleSense test set is present in the combined train, dev, and test datasets of BrainTeaser. To achieve this, we are using the Sentence Transformers library  [^5] SentenceTransformers Reimers and Gurevych (2019) with the gte-large-en-v1.5  [^6] Alibaba-NLP/gte-large-en-v1.5 Zhang et al. (2024) model to generate text embeddings and perform semantic similarity comparisons. We convert each question from the entire BrainTeaser puzzle into embeddings that represent the semantic meaning of each question. We then calculate the cosine similarity between these embeddings and those in the RiddleSense test set to identify duplicates or highly similar questions. A lower threshold is applied to define similarity, meaning that any question with a cosine similarity above this threshold is considered a near-duplicate and excluded from the test set. After conducting several experiments, we determined that a cosine similarity threshold of 0.9 was optimal. This threshold is a delicate balance, as our questions are riddle-style and often rely on abstract or metaphorical language, making it challenging to clearly assess whether two questions are similar enough to warrant exclusion. Since riddles are not always straightforward in their wording, the threshold needed to be fine-tuned to avoid excluding questions that are conceptually distinct but might have overlapping language, ensuring a fair evaluation without overly aggressive filtering. This means that any question from the RiddleSense test set that had a similarity score of 0.9 or higher to a question in the BrainTeaser datasets was removed to avoid redundancy. As previously mentioned, because we prioritize BrainTeaser questions due to their superior quality from manual annotation, we decided to remove any similar questions from the RiddleSense test set. This approach helps preserve the integrity of the evaluation process and ensures that the test set contains only unique and distinct questions, minimizing potential data leakage and improving the robustness of our results. Additionally, by removing these instances, we can perform more experiments on the remaining unique questions, optimizing our use of hardware resources and potentially yielding more comprehensive insights.
Appendix E Prompting details
Zero-shot Chain of Thought Prompting Technique
This is the most straightforward prompting technique. In this approach, each instance from the test set of each dataset is presented to the model using a specific system-user prompt format. This technique operates in a zero-shot setting, meaning the model is not given any prior examples or training specific to the task at hand. Instead, it relies solely on the system prompt and user prompt to generate responses based on its pre-existing knowledge. The chaoin of thought character is a result of the usage of the phrase Let’s think step by step in the end of each system prompt. The Chain of Thought (CoT) character is achieved by including the phrase “Let’s think step by step” at the end of each system prompt. This approach aligns with the findings of Kojima et al. (2023), which demonstrate that large language models can perform effectively as zero-shot reasoners by incorporating this specific phrase before presenting an unanswered test instance. The following are the system-user prompts that were used:
  System Prompt:
You will encounter a riddle that requires analytical thinking and reasoning to solve.
A riddle is a question or statement intentionally phrased so as to require ingenuity in ascertaining its answer or meaning, typically presented as a game.
Different ideas can be used in these riddles:
1. Riddles often employ misdirection, leading you away from the actual solution.
2. They include elements with double meanings, requiring a keen eye for words with dual interpretations.
3. Metaphorical wordplay adds another layer, urging you to decipher figurative language.
4. Look out for exaggeration, as riddles may present overly dramatic details to divert your attention.
5. Common phrases and sayings may hide within the puzzle, demanding familiarity.
6. Associations and irony play a crucial role, introducing unexpected connections.
7. Numerical puzzles can also be part of the mystery, requiring you to decode their significance.
8. Elemental imagery, drawn from nature, might hold key descriptors.
9. Rhyming and sound clues can add a poetic dimension. 10. Word Puzzles: Pay attention to anagrams, acrostics, and other wordplay elements.
11. Also, it is important to note you should decode the upcoming riddle using everyday logic and creativity.
Approach the riddle with everyday logic and creativity, avoiding supernatural explanations.
You will be given an unsolved riddle and five options to choose the answer amongst them. Let’s think step by step.
User Prompt:
Riddle: ‘‘‘
{RIDDLE}
Options:
[option 1]: ‘‘‘{OPTION_1}‘‘‘
[option 2]: ‘‘‘{OPTION_2}‘‘‘
[option 3]: ‘‘‘{OPTION_3}‘‘‘
[option 4]: ‘‘‘{OPTION_4}‘‘‘
[option 5]: ‘‘‘{OPTION_5}‘‘‘
  Where the input parameters are self-explanatory. It is worth mentioning that for BrainTeaser, where the multiple-choice options are four, the prompt is adjusted accordingly.
Few-shot Prompting Techniques
In this approach, we provide a specific number of exemplars from each dataset’s training set before asking the model to solve a multiple-choice question from the test set. These exemplars, which include their correct answers, serve as in-context learning examples. By presenting these examples, we aim to guide the model and improve its performance on the subsequent unanswered tasks. This method leverages the provided context to enhance the model’s ability to understand and solve new instances more effectively. The system prompt that is used here is the one of the zero-shot prompting with some minor changes. The number of examples that are used are 2, 4 and 8 as mentioned in section 3.
Now, a key decision is how to select the exemplars that will be provided to the model before it attempts to answer the test set questions. The simplest and most straightforward approach is to randomly choose these answered exemplars from the training dataset. This method, referred to as FS Rand in our experiments, offers a baseline for comparison, as it doesn’t apply any specific strategy for selecting the most relevant examples and relies purely on random sampling.
Another approach implemented in our experiments for selecting exemplars is based on semantic similarity. Specifically, we utilize the Sentence Transformers library  [^7] SentenceTransformers Reimers and Gurevych (2019) in conjunction with the gte-large-en-v1.5 model  [^8] Alibaba-NLP/gte-large-en-v1.5 Zhang et al. (2024). This setup allows us to generate text embeddings and perform semantic similarity comparisons using cosine similarity between these embeddings. For each experiment, we first generate text embeddings for all instances in the training datasets. Then, for each question in the test set, we generate its embedding and calculate its cosine similarity with all training set embeddings. Based on the number of in-context learning examples needed, we select the most similar training instances to use as exemplars. This method ensures that the selected examples are semantically relevant, aiming to improve the model’s performance by presenting it with contextually aligned training examples. This approach is referred to as FS Sim in our experiments, as it focuses on semantic similarity to guide exemplar selection for improved task performance.
The system and user prompts used for both of the above settings are outlined as follows:
  System Prompt:
You will encounter a riddle that requires analytical thinking and reasoning to solve.
A riddle is a question or statement intentionally phrased so as to require ingenuity in ascertaining its answer or meaning, typically presented as a game.
Different ideas can be used in these riddles:
1. Riddles often employ misdirection, leading you away from the actual solution.
2. They include elements with double meanings, requiring a keen eye for words with dual interpretations.
3. Metaphorical wordplay adds another layer, urging you to decipher figurative language.
4. Look out for exaggeration, as riddles may present overly dramatic details to divert your attention.
5. Common phrases and sayings may hide within the puzzle, demanding familiarity.
6. Associations and irony play a crucial role, introducing unexpected connections.
7. Numerical puzzles can also be part of the mystery, requiring you to decode their significance.
8. Elemental imagery, drawn from nature, might hold key descriptors.
9. Rhyming and sound clues can add a poetic dimension.
10. Word Puzzles: Pay attention to anagrams, acrostics, and other wordplay elements.
11. Also, it is important to note you should decode the upcoming riddle using everyday logic and creativity.
Approach the riddle with everyday logic and creativity, avoiding supernatural explanations.
First, you’ll encounter X examples with their answer provided similar to the riddle you will need to solve.
Then you will be given an unsolved riddle and X options to choose the answer amongst them.
User Prompt:
{EXAMPLES}
Answer the following riddle while taking into consideration the examples above. Choose the best and the most logical option from the available choices:
Riddle: ‘‘‘
{RIDDLE}
Options:
[option 1]: ‘‘‘{OPTION_1}‘‘‘
[option 2]: ‘‘‘{OPTION_2}‘‘‘
[option 3]: ‘‘‘{OPTION_3}‘‘‘
[option 4]: ‘‘‘{OPTION_4}‘‘‘
[option 5]: ‘‘‘{OPTION_5}‘‘‘
  Where the EXAMPLES refer to the few-shot examples selected for each test set instance, and the other input parameters are self-explanatory. It is worth mentioning that for BrainTeaser, where the multiple-choice options are four, the prompt is adjusted accordingly.
Chain of Thought Prompting Technique
Another promising technique for enhancing model performance across various tasks is Chain of Thought (CoT) prompting. This approach involves not only providing answered examples but also including an explanation consisting of intermediate reasoning steps. By laying out the thought process step-by-step, CoT prompting significantly improves the capability of large language models to tackle complex reasoning tasks. In our task, the Chain of Thought (CoT) prompting technique is anticipated to outperform the previously mentioned methods. This expectation is supported by the performance improvements observed in the CommonsenseQA dataset Talmor et al. (2019), a domain similar to our setting that involves both lateral and vertical thinking. The dataset demonstrated significant improvements when using a manually composed set of two, four, or eight few-shot exemplars with CoT prompting, highlighting its effectiveness in eliciting successful reasoning. In our case, these explanations were generated manually to ensure they align with human perceptions of the reasoning process required by the riddle. The explanation format we used follows the structure outlined by Wei et al. (2023) for few-shot exemplars in full chain-of-thought prompts for CommonsenseQA. Due to the similarity in the multiple-choice question-answer format of our datasets, it served as a good foundation for developing our own approach of creating explanations. Below, we provide the system and user prompts used in the aforementioned technique:
  System Prompt:
You will encounter a riddle that requires analytical thinking and reasoning to solve.
A riddle is a question or statement intentionally phrased so as to require ingenuity in ascertaining its answer or meaning, typically presented as a game.
Different ideas can be used in these riddles:
1. Riddles often employ misdirection, leading you away from the actual solution.
2. They include elements with double meanings, requiring a keen eye for words with dual interpretations.
3. Metaphorical wordplay adds another layer, urging you to decipher figurative language.
4. Look out for exaggeration, as riddles may present overly dramatic details to divert your attention.
5. Common phrases and sayings may hide within the puzzle, demanding familiarity.
6. Associations and irony play a crucial role, introducing unexpected connections.
7. Numerical puzzles can also be part of the mystery, requiring you to decode their significance.
8. Elemental imagery, drawn from nature, might hold key descriptors.
9. Rhyming and sound clues can add a poetic dimension.
10. Word Puzzles: Pay attention to anagrams, acrostics, and other wordplay elements.
11. Also, it is important to note you should decode the upcoming riddle using everyday logic and creativity.
Approach the riddle with everyday logic and creativity, avoiding supernatural explanations.
First, you’ll encounter X examples demonstrating analytical reasoning similar to the riddle you will need to solve.
Then you will be given an unsolved riddle and five options to choose the answer amongst them. Let’s think step by step and solve the riddle based on the examples provided above.
User Prompt:
{EXAMPLES_COT}
Answer the following riddle while taking into consideration the examples above. Choose the best and the most logical option from the available choices:
Riddle: ‘‘‘
{RIDDLE}
Options:
[option 1]: ‘‘‘{OPTION_1}‘‘‘
[option 2]: ‘‘‘{OPTION_2}‘‘‘
[option 3]: ‘‘‘{OPTION_3}‘‘‘
[option 4]: ‘‘‘{OPTION_4}‘‘‘
[option 5]: ‘‘‘{OPTION_5}‘‘‘
  Where the EXAMPLES_COT refer to the few-shot examples with manually generated explanations selected for each test set instance, and the other input parameters are self-explanatory. It is important to note that for BrainTeaser, where there are four multiple-choice options, the prompt is modified accordingly.
Appendix F Experimental Setup
In our experiments, we employed the Google Colab platform and Kaggle, leveraging various open-source Python packages such as Transformers, BitsAndBytes, Accelerate Gugger et al. (2022) and Sentence-Transformers. We also utilized the Amazon Bedrock API, which allowed us to access models without being constrained by hardware limitations, providing scalable and flexible model deployment.
We specifically opted for instruction variations of the models because they aligned more closely with our task requirements. Instruction-tuned models generally offer better performance in scenarios where understanding and following specific instructions is crucial. This alignment ensures that the models are better equipped to handle the nuances of the tasks, leading to more effective and relevant outputs.
The temperature and repetition penalty values were determined through a series of exploratory experiments. To ensure consistency across our work, we systematically applied the same parameters whenever possible. For our experiments, we used a temperature of 0.5 and repetition penalties of either 1.0 or 1.15.
Llama 3
AI@Meta (2024) In our experiments, we chose two variations of the model: the 8B and 70B versions. The Llama3-8B999meta-llama/Meta-Llama-3-8B-Instruct model was used for inference without quantization, which allowed it to deliver results with full precision. This approach was ideal when computational resources were sufficient and precision was critical. On the other hand, the Llama3-70B 101010meta-llama/Meta-Llama-3-70B-Instruct model, due to hardware limitations, underwent quantization. This process reduced the model’s size and computational needs, making it accessible despite the constraints of our hardware. While quantization might lead to some loss of precision, it was necessary to deploy a model with more parameters, which otherwise would not have been feasible.
Mistral
Jiang et al. (2023a) In our experiments, we used the Mistral-7B-Instruct-v0.2111111mistralai/Mistral-7B-Instruct-v0.2 and the Mixtral-8x7B-Instruct-v0.1121212mistralai/Mixtral-8x7B-Instruct-v0.1 in their unquantized forms. The Mistral 7B-Instruct v0.2 was selected for its strong instruction-following abilities, maintaining full precision. Similarly, the Mixtral-8x7B-Instruct-v0.1, which integrates eight 7B models, was used unquantized to benefit from its combined performance. Our goal was to explore the effectiveness of smaller models for our riddle tasks, and we preferred these instruction-tuned variations to ensure they were well-suited to our specific requirements, leading to more effective results.
Qwen2
Yang et al. (2024) To further explore smaller model variations, we selected the Qwen2-7B-Instruct [^13] Qwen/Qwen2-7B-Instruct version. The Qwen2 family of models, including this 7B-Instruct variant, is known for its strong performance in reasoning tasks. The Qwen2 models are designed with advanced instruction-following capabilities and are particularly effective at complex problem-solving and logical reasoning, making them well-suited for our objectives.
Appendix G Detailed Results
In this section, we present the detailed results for both datasets across all our experimental techniques. Due to the extensive nature of the experiments, the results for each dataset are organized into subtables. We begin with the BrainTeaser dataset, where the metrics are more detailed due to the dataset’s structure, followed by the results for the RiddleSense dataset. The structure is organized by method per model, with results presented in descending order based on score for each method, rather than by the number of examples. This differs from the approach outlined in 5.
Model Method Num.Ex Task Temp Rep_Pen Original Semantic Context Ori. + Sem. Ori. + Sem. + Con. Average
Chain-of-Thought Zero-shot
Meta-Llama-3-70B-Instruct CoT_ZS (Q) 0 SP 0.5 1.15 0.725 0.775 0.675 0.675 0.550 0.725
Mixtral-8x7B-Instruct-v0.1 CoT_ZS 0 SP 0.5 1.0 0.575 0.550 0.525 0.475 0.275 0.550
Meta-Llama-3-8B-Instruct CoT_ZS 0 SP 0.5 1.0 0.625 0.650 0.625 0.500 0.325 0.633
Mistral-7B-Instruct-v0.2 CoT_ZS 0 SP 0.5 1.0 0.375 0.475 0.500 0.300 0.250 0.450
Qwen2-7B-Instruct CoT_ZS 0 SP 0.5 1.15 0.475 0.450 0.450 0.350 0.200 0.458
Few-shot with CoT Explanations
Meta-Llama-3-70B-Instruct CoT_FS (Q) 2 SP 0.5 1.15 0.850 0.750 0.675 0.700 0.500 0.758
Meta-Llama-3-70B-Instruct CoT_FS (Q) 8 SP 0.5 1.0 0.675 0.750 0.700 0.600 0.425 0.708
Meta-Llama-3-70B-Instruct CoT_FS (Q) 4 SP 0.5 1.15 0.700 0.650 0.700 0.575 0.450 0.683
Mixtral-8x7B-v0.1 CoT_FS 8 SP 0.5 1.0 0.650 0.675 0.600 0.575 0.375 0.642
Mixtral-8x7B-v0.1 CoT_FS 2 SP 0.5 1.0 0.625 0.625 0.600 0.525 0.350 0.617
Mixtral-8x7B-v0.1 CoT_FS 4 SP 0.5 1.0 0.575 0.600 0.575 0.475 0.350 0.583
Meta-Llama-3-8B-Instruct CoT_FS 8 SP 0.5 1.0 0.675 0.700 0.600 0.525 0.325 0.658
Meta-Llama-3-8B-Instruct CoT_FS 2 SP 0.5 1.0 0.725 0.625 0.550 0.550 0.350 0.633
Meta-Llama-3-8B-Instruct CoT_FS 4 SP 0.5 1.0 0.650 0.625 0.550 0.500 0.300 0.608
Mistral-7B-Instruct-v0.2 CoT_FS 8 SP 0.5 1.0 0.525 0.550 0.450 0.375 0.300 0.508
Mistral-7B-Instruct-v0.2 CoT_FS 4 SP 0.5 1.0 0.525 0.550 0.450 0.375 0.300 0.508
Mistral-7B-Instruct-v0.2 CoT_FS 2 SP 0.5 1.0 0.525 0.425 0.475 0.300 0.225 0.475
Qwen2-7B-Instruct CoT_FS 8 SP 0.5 1.0 0.600 0.725 0.675 0.550 0.425 0.667
Qwen2-7B-Instruct CoT_FS 4 SP 0.5 1.0 0.650 0.675 0.625 0.550 0.450 0.650
Qwen2-7B-Instruct CoT_FS 2 SP 0.5 1.0 0.675 0.550 0.600 0.450 0.375 0.608
Few-shot with Random Selection
Meta-Llama-3-70B-Instruct FS Rand (Q) 4 SP 0.5 1.15 0.825 0.850 0.750 0.800 0.700 0.808
Meta-Llama-3-70B-Instruct FS Rand (Q) 8 SP 0.5 1.15 0.775 0.800 0.750 0.675 0.550 0.775
Meta-Llama-3-70B-Instruct FS Rand (Q) 2 SP 0.5 1.15 0.750 0.800 0.775 0.725 0.600 0.775
Mixtral-8x7B-Instruct-v0.1 FS Rand 4 SP 0.5 1.0 0.675 0.775 0.600 0.600 0.400 0.683
Mixtral-8x7B-Instruct-v0.1 FS Rand 2 SP 0.5 1.0 0.650 0.625 0.575 0.475 0.300 0.617
Mixtral-8x7B-Instruct-v0.1 FS Rand 8 SP 0.5 1.0 0.675 0.650 0.525 0.525 0.350 0.617
Meta-Llama-3-8B-Instruct FS Rand 8 SP 0.5 1.0 0.750 0.625 0.650 0.575 0.400 0.675
Meta-Llama-3-8B-Instruct FS Rand 4 SP 0.5 1.0 0.625 0.650 0.650 0.500 0.375 0.642
Meta-Llama-3-8B-Instruct FS Rand 2 SP 0.5 1.0 0.600 0.700 0.600 0.550 0.375 0.633
Mistral-7B-Instruct-v0.2 FS Rand 2 SP 0.5 1.0 0.525 0.550 0.475 0.425 0.275 0.517
Mistral-7B-Instruct-v0.2 FS Rand 8 SP 0.5 1.0 0.450 0.575 0.425 0.400 0.300 0.483
Mistral-7B-Instruct-v0.2 FS Rand 4 SP 0.5 1.0 0.550 0.425 0.475 0.325 0.225 0.483
Qwen2-7B-Instruct FS Rand 2 SP 0.5 1.15 0.675 0.650 0.600 0.550 0.425 0.642
Qwen2-7B-Instruct FS Rand 8 SP 0.5 1.15 0.700 0.600 0.625 0.525 0.425 0.642
Qwen2-7B-Instruct FS Rand 4 SP 0.5 1.15 0.675 0.575 0.575 0.525 0.400 0.608
Table 6: Model performance for BrainTeaser (Part 1). The presence of (Q) in the method column indicates that the results correspond to the quantized version of the model.
Model Method Num.Ex Task Temp Rep_Pen Original Semantic Context Ori. + Sem. Ori. + Sem. + Con. Average
Few-shot with Semantic Similarity
Meta-Llama-3-70B-Instruct FS Sim (Q) 2 SP 0.5 1.15 0.850 0.875 0.750 0.800 0.675 0.825
Meta-Llama-3-70B-Instruct FS Sim (Q) 4 SP 0.5 1.15 0.825 0.775 0.775 0.700 0.575 0.792
Meta-Llama-3-70B-Instruct FS Sim (Q) 8 SP 0.5 1.15 0.775 0.825 0.750 0.700 0.600 0.783
Mixtral-8x7B-Instruct-v0.1 FS Sim 2 SP 0.5 1.0 0.700 0.700 0.675 0.575 0.450 0.692
Mixtral-8x7B-Instruct-v0.1 FS Sim 4 SP 0.5 1.0 0.750 0.650 0.650 0.575 0.425 0.683
Mixtral-8x7B-Instruct-v0.1 FS Sim 8 SP 0.5 1.0 0.750 0.625 0.625 0.525 0.375 0.667
Meta-Llama-3-8B-Instruct FS Sim 8 SP 0.5 1.0 0.750 0.775 0.775 0.700 0.600 0.767
Meta-Llama-3-8B-Instruct FS Sim 4 SP 0.5 1.0 0.675 0.825 0.650 0.675 0.475 0.717
Meta-Llama-3-8B-Instruct FS Sim 2 SP 0.5 1.0 0.800 0.675 0.625 0.625 0.525 0.700
Mistral-7B-Instruct-v0.2 FS Sim 8 SP 0.5 1.0 0.550 0.600 0.450 0.500 0.325 0.533
Mistral-7B-Instruct-v0.2 FS Sim 2 SP 0.5 1.0 0.55 0.475 0.525 0.375 0.275 0.517
Mistral-7B-Instruct-v0.2 FS Sim 4 SP 0.5 1.0 0.475 0.450 0.450 0.350 0.225 0.458
Qwen2-7B-Instruct FS Sim 8 SP 0.5 1.15 0.675 0.625 0.650 0.575 0.500 0.650
Qwen2-7B-Instruct FS Sim 4 SP 0.5 1.15 0.650 0.600 0.650 0.550 0.450 0.633
Qwen2-7B-Instruct FS Sim 2 SP 0.5 1.15 0.600 0.550 0.650 0.500 0.400 0.600
RISCOREm Rand
Meta-Llama-3-70B-Instruct RISCOREm Rand (Q) 8 SP 0.5 1.15 0.775 0.825 0.800 0.700 0.600 0.800
Meta-Llama-3-70B-Instruct RISCOREm Rand (Q) 2 SP 0.5 1.15 0.850 0.775 0.725 0.725 0.650 0.783
Meta-Llama-3-70B-Instruct RISCOREm Rand (Q) 4 SP 0.5 1.15 0.775 0.800 0.700 0.725 0.600 0.758
Mixtral-8x7B-Instruct-v0.1 RISCOREm Rand 8 SP 0.5 1.0 0.700 0.625 0.625 0.600 0.450 0.650
Mixtral-8x7B-Instruct-v0.1 RISCOREm Rand 2 SP 0.5 1.0 0.725 0.600 0.550 0.600 0.400 0.625
Mixtral-8x7B-Instruct-v0.1 RISCOREm Rand 4 SP 0.5 1.0 0.675 0.650 0.525 0.550 0.325 0.617
Meta-Llama-3-8B-Instruct RISCOREm Rand 4 SP 0.5 1.0 0.650 0.700 0.675 0.625 0.500 0.675
Meta-Llama-3-8B-Instruct RISCOREm Rand 2 SP 0.5 1.0 0.700 0.725 0.575 0.525 0.375 0.667
Meta-Llama-3-8B-Instruct RISCOREm Rand 8 SP 0.5 1.0 0.700 0.750 0.550 0.575 0.425 0.667
Mistral-7B-Instruct-v0.2 RISCOREm Rand 4 SP 0.5 1.0 0.575 0.475 0.500 0.425 0.300 0.517
Mistral-7B-Instruct-v0.2 RISCOREm Rand 2 SP 0.5 1.0 0.550 0.425 0.400 0.325 0.200 0.458
Mistral-7B-Instruct-v0.2 RISCOREm Rand 8 SP 0.5 1.0 0.350 0.450 0.400 0.275 0.200 0.400
Qwen2-7B-Instruct RISCOREm Rand 4 SP 0.5 1.15 0.625 0.650 0.600 0.550 0.400 0.625
Qwen2-7B-Instruct RISCOREm Rand 2 SP 0.5 1.15 0.650 0.575 0.600 0.500 0.400 0.608
Qwen2-7B-Instruct RISCOREm Rand 8 SP 0.5 1.15 0.625 0.550 0.600 0.450 0.325 0.592
RISCOREm Sim
Meta-Llama-3-70B-Instruct RISCOREm Sim (Q) 4 SP 0.5 1.15 0.850 0.850 0.800 0.800 0.650 0.833
Meta-Llama-3-70B-Instruct RISCOREm Sim (Q) 8 SP 0.5 1.15 0.800 0.775 0.850 0.700 0.625 0.808
Meta-Llama-3-70B-Instruct RISCOREm Sim (Q) 2 SP 0.5 1.15 0.850 0.775 0.725 0.725 0.600 0.783
Mixtral-8x7B-Instruct-v0.1 RISCOREm Sim 8 SP 0.5 1.0 0.775 0.625 0.725 0.575 0.500 0.708
Mixtral-8x7B-Instruct-v0.1 RISCOREm Sim 4 SP 0.5 1.0 0.675 0.750 0.700 0.600 0.425 0.708
Mixtral-8x7B-Instruct-v0.1 RISCOREm Sim 2 SP 0.5 1.0 0.725 0.675 0.625 0.675 0.475 0.675
Meta-Llama-3-8B-Instruct RISCOREm Sim 2 SP 0.5 1.0 0.750 0.875 0.675 0.675 0.575 0.767
Meta-Llama-3-8B-Instruct RISCOREm Sim 8 SP 0.5 1.0 0.775 0.800 0.700 0.775 0.650 0.758
Meta-Llama-3-8B-Instruct RISCOREm Sim 4 SP 0.5 1.0 0.825 0.700 0.700 0.675 0.525 0.742
Mistral-7B-Instruct-v0.2 RISCOREm Sim 4 SP 0.5 1.0 0.575 0.550 0.575 0.400 0.325 0.567
Mistral-7B-Instruct-v0.2 RISCOREm Sim 8 SP 0.5 1.0 0.575 0.575 0.500 0.450 0.325 0.550
Mistral-7B-Instruct-v0.2 RISCOREm Sim 2 SP 0.5 1.0 0.500 0.475 0.475 0.350 0.200 0.483
Qwen2-7B-Instruct RISCOREm Sim 8 SP 0.5 1.15 0.700 0.650 0.650 0.575 0.425 0.667
Qwen2-7B-Instruct RISCOREm Sim 2 SP 0.5 1.15 0.700 0.625 0.675 0.550 0.375 0.667
Qwen2-7B-Instruct RISCOREm Sim 4 SP 0.5 1.15 0.750 0.625 0.550 0.575 0.425 0.642
Table 7: Model performance for BrainTeaser (Part 2). The presence of (Q) in the method column indicates that the results correspond to the quantized version of the model.
Model Method Num.Ex Task Temp Rep_Pen Original Semantic Context Ori. + Sem. Ori. + Sem. + Con. Average
RISCORE Results
Llama3-70B zeroshot for QA & Llama3-8B for distractors
Meta-Llama-3-70B-Instruct RISCORE Sim (Q) 8 SP 0.5 1.15 0.850 0.825 0.750 0.775 0.600 0.808
Meta-Llama-3-70B-Instruct RISCORE Sim (Q) 2 SP 0.5 1.15 0.800 0.775 0.800 0.750 0.650 0.792
Meta-Llama-3-70B-Instruct RISCORE Sim (Q) 4 SP 0.5 1.15 0.825 0.750 0.800 0.675 0.600 0.792
Mixtral-8x7B-Instruct-v0.1 RISCORE Sim 8 SP 0.5 1.0 0.700 0.750 0.600 0.650 0.475 0.683
Mixtral-8x7B-Instruct-v0.1 RISCORE Sim 2 SP 0.5 1.0 0.750 0.750 0.500 0.650 0.375 0.667
Mixtral-8x7B-Instruct-v0.1 RISCORE Sim 4 SP 0.5 1.0 0.725 0.625 0.575 0.575 0.375 0.642
Meta-Llama-3-8B-Instruct RISCORE Sim 8 SP 0.5 1.0 0.700 0.725 0.675 0.650 0.525 0.700
Meta-Llama-3-8B-Instruct RISCORE Sim 4 SP 0.5 1.0 0.650 0.675 0.700 0.550 0.425 0.675
Meta-Llama-3-8B-Instruct RISCORE Sim 2 SP 0.5 1.0 0.700 0.600 0.575 0.525 0.350 0.625
Mistral-7B-Instruct-v0.2 RISCORE Sim 2 SP 0.5 1.0 0.550 0.450 0.475 0.375 0.300 0.492
Mistral-7B-Instruct-v0.2 RISCORE Sim 4 SP 0.5 1.0 0.525 0.400 0.475 0.300 0.200 0.467
Mistral-7B-Instruct-v0.2 RISCORE Sim 8 SP 0.5 1.0 0.425 0.525 0.475 0.400 0.325 0.475
Qwen2-7B-Instruct RISCORE Sim 8 SP 0.5 1.15 0.600 0.700 0.625 0.550 0.425 0.642
Qwen2-7B-Instruct RISCORE Sim 2 SP 0.5 1.15 0.650 0.650 0.575 0.550 0.400 0.625
Qwen2-7B-Instruct RISCORE Sim 4 SP 0.5 1.15 0.600 0.650 0.625 0.550 0.450 0.625
Llama3-70B fewshot for QA & Llama3-8B for distractors
Meta-Llama-3-70B-Instruct RISCORE Sim (Q) 8 SP 0.5 1.15 0.800 0.800 0.825 0.750 0.650 0.808
Meta-Llama-3-70B-Instruct RISCORE Sim (Q) 4 SP 0.5 1.15 0.775 0.800 0.800 0.700 0.600 0.792
Meta-Llama-3-70B-Instruct RISCORE Sim (Q) 2 SP 0.5 1.15 0.800 0.725 0.725 0.725 0.600 0.750
Mixtral-8x7B-Instruct-v0.1 RISCORE Sim 2 SP 0.5 1.0 0.725 0.725 0.575 0.675 0.475 0.675
Mixtral-8x7B-Instruct-v0.1 RISCORE Sim 8 SP 0.5 1.0 0.700 0.675 0.65 0.625 0.450 0.675
Mixtral-8x7B-Instruct-v0.1 RISCORE Sim 4 SP 0.5 1.0 0.700 0.675 0.575 0.625 0.400 0.650
Meta-Llama-3-8B-Instruct RISCORE Sim 8 SP 0.5 1.0 0.775 0.750 0.700 0.700 0.600 0.742
Meta-Llama-3-8B-Instruct RISCORE Sim 2 SP 0.5 1.0 0.725 0.700 0.625 0.600 0.400 0.683
Meta-Llama-3-8B-Instruct RISCORE Sim 4 SP 0.5 1.0 0.725 0.625 0.625 0.550 0.450 0.658
Mistral-7B-Instruct-v0.2 RISCORE Sim 4 SP 0.5 1.0 0.625 0.575 0.475 0.475 0.350 0.558
Mistral-7B-Instruct-v0.2 RISCORE Sim 8 SP 0.5 1.0 0.500 0.550 0.500 0.425 0.350 0.517
Mistral-7B-Instruct-v0.2 RISCORE Sim 2 SP 0.5 1.0 0.550 0.425 0.450 0.375 0.225 0.475
Qwen2-7B-Instruct RISCORE Sim 4 SP 0.5 1.15 0.625 0.700 0.650 0.575 0.450 0.658
Qwen2-7B-Instruct RISCORE Sim 8 SP 0.5 1.15 0.650 0.675 0.650 0.575 0.475 0.658
Qwen2-7B-Instruct RISCORE Sim 2 SP 0.5 1.15 0.625 0.650 0.600 0.525 0.400 0.625
Llama3-70B fewshot for QA & Llama3-70B for distractors
Meta-Llama-3-70B-Instruct RISCORE Sim (Q) 4 SP 0.5 1.15 0.875 0.775 0.725 0.750 0.600 0.792
Meta-Llama-3-70B-Instruct RISCORE Sim (Q) 2 SP 0.5 1.15 0.775 0.825 0.750 0.775 0.675 0.783
Meta-Llama-3-70B-Instruct RISCORE Sim (Q) 8 SP 0.5 1.15 0.775 0.750 0.775 0.725 0.600 0.767
Mixtral-8x7B-Instruct-v0.1 RISCORE Sim 8 SP 0.5 1.0 0.700 0.725 0.625 0.600 0.450 0.683
Mixtral-8x7B-Instruct-v0.1 RISCORE Sim 2 SP 0.5 1.0 0.700 0.700 0.600 0.625 0.500 0.667
Mixtral-8x7B-Instruct-v0.1 RISCORE Simm 4 SP 0.5 1.0 0.725 0.650 0.550 0.600 0.425 0.642
Meta-Llama-3-8B-Instruct RISCORE Sim 8 SP 0.5 1.0 0.800 0.675 0.625 0.625 0.475 0.700
Meta-Llama-3-8B-Instruct RISCORE Sim 2 SP 0.5 1.0 0.675 0.700 0.675 0.600 0.475 0.683
Meta-Llama-3-8B-Instruct RISCORE Sim 4 SP 0.5 1.0 0.725 0.650 0.625 0.550 0.475 0.667
Mistral-7B-Instruct-v0.2 RISCORE Sim 4 SP 0.5 1.0 0.575 0.500 0.450 0.375 0.300 0.508
Mistral-7B-Instruct-v0.2 RISCORE Sim 2 SP 0.5 1.0 0.625 0.400 0.475 0.350 0.275 0.500
Mistral-7B-Instruct-v0.2 RISCORE Sim 8 SP 0.5 1.0 0.600 0.475 0.425 0.400 0.300 0.500
Qwen2-7B-Instruct RISCORE Sim 2 SP 0.5 1.15 0.650 0.600 0.600 0.500 0.400 0.617
Qwen2-7B-Instruct RISCORE Sim 4 SP 0.5 1.15 0.650 0.625 0.575 0.600 0.475 0.617
Qwen2-7B-Instruct RISCORE Sim 8 SP 0.5 1.15 0.625 0.625 0.600 0.575 0.450 0.617
Table 8: Model performance for BrainTeaser (Part 3). The presence of (Q) in the method column indicates that the results correspond to the quantized version of the model.
Model Method Num.Ex Quant Temp Rep_Pen Average
Chain-of-Thought Zero-shot
Meta-Llama-3-70B-Instruct CoT_ZS (Q) 0 4bit 0.5 1.15 0.775
Mixtral-8x7B-v0.1 CoT_ZS 0 False 0.5 1.0 0.675
Meta-Llama-3-8B-Instruct CoT_ZS 0 False 0.5 1.0 0.619
Mistral-7B-Instruct-v0.2 CoT_ZS 0 False 0.5 1.0 0.589
Qwen2-7B-Instruct CoT_ZS 0 False 0.5 1.15 0.608
Meta-Llama-3-8B-Instruct CoT_ZS 0 4bit 0.5 1.15 0.633
Few-shot with CoT Explanations
Meta-Llama-3-70B-Instruct CoT_FS (Q) 2 4bit 0.5 1.15 0.789
Meta-Llama-3-70B-Instruct CoT_FS (Q) 4 4bit 0.5 1.15 0.783
Meta-Llama-3-70B-Instruct CoT_FS (Q) 8 4bit 0.5 1.15 0.783
Mixtral-8x7B-v0.1 CoT_FS 8 False 0.5 1.0 0.697
Mixtral-8x7B-v0.1 CoT_FS 2 False 0.5 1.0 0.692
Mixtral-8x7B-v0.1 CoT_FS 4 False 0.5 1.0 0.686
Meta-Llama-3-8B-Instruct CoT_FS 4 False 0.5 1.0 0.672
Meta-Llama-3-8B-Instruct CoT_FS 8 False 0.5 1.0 0.658
Meta-Llama-3-8B-Instruct CoT_FS 2 False 0.5 1.0 0.625
Mistral-7B-Instruct-v0.2 CoT_FS 4 False 0.5 1.0 0.603
Mistral-7B-Instruct-v0.2 CoT_FS 8 False 0.5 1.0 0.597
Mistral-7B-Instruct-v0.2 CoT_FS 2 False 0.5 1.0 0.594
Qwen2-7B-Instruct CoT_FS 2 False 0.5 1.15 0.667
Qwen2-7B-Instruct CoT_FS 4 False 0.5 1.15 0.656
Qwen2-7B-Instruct CoT_FS 8 False 0.5 1.15 0.625
Few-shot with Random Selection
Meta-Llama-3-70B-Instruct FS Rand (Q) 4 4bit 0.5 1.15 0.800
Meta-Llama-3-70B-Instruct FS Rand (Q) 8 4bit 0.5 1.15 0.772
Meta-Llama-3-70B-Instruct FS Rand (Q) 2 4bit 0.5 1.15 0.769
Mixtral-8x7B-v0.1 FS Rand 4 False 0.5 1.0 0.719
Mixtral-8x7B-v0.1 FS Rand 8 False 0.5 1.0 0.711
Mixtral-8x7B-v0.1 FS Rand 2 False 0.5 1.0 0.706
Meta-Llama-3-8B-Instruct FS Rand 2 False 0.5 1.0 0.672
Meta-Llama-3-8B-Instruct FS Rand 8 False 0.5 1.0 0.672
Meta-Llama-3-8B-Instruct FS Rand 4 False 0.5 1.0 0.639
Mistral-7B-Instruct-v0.2 FS Rand 2 False 0.5 1.0 0.586
Mistral-7B-Instruct-v0.2 FS Rand 4 False 0.5 1.0 0.586
Mistral-7B-Instruct-v0.2 FS Rand 8 False 0.5 1.0 0.586
Qwen2-7B-Instruct FS Rand 8 False 0.5 1.15 0.700
Qwen2-7B-Instruct FS Rand 2 False 0.5 1.15 0.689
Qwen2-7B-Instruct FS Rand 4 False 0.5 1.15 0.683
Few-shot with Semantic Similarity
Meta-Llama-3-70B-Instruct FS Sim (Q) 4 4bit 0.5 1.15 0.817
Meta-Llama-3-70B-Instruct FS Sim (Q) 8 4bit 0.5 1.15 0.800
Meta-Llama-3-70B-Instruct FS Sim (Q) 2 4bit 0.5 1.15 0.792
Mixtral-8x7B-v0.1 FS Sim 2 False 0.5 1.0 0.714
Mixtral-8x7B-v0.1 FS Sim 4 False 0.5 1.0 0.692
Mixtral-8x7B-v0.1 FS Sim 8 False 0.5 1.0 0.675
Meta-Llama-3-8B-Instruct FS Sim 4 False 0.5 1.0 0.711
Meta-Llama-3-8B-Instruct FS Sim 2 False 0.5 1.0 0.706
Meta-Llama-3-8B-Instruct FS Sim 8 False 0.5 1.0 0.681
Mistral-7B-Instruct-v0.2 FS Sim 4 False 0.5 1.0 0.633
Mistral-7B-Instruct-v0.2 FS Sim 8 False 0.5 1.0 0.611
Mistral-7B-Instruct-v0.2 FS Sim 2 False 0.5 1.0 0.608
Qwen2-7B-Instruct FS Sim 8 False 0.5 1.15 0.731
Qwen2-7B-Instruct FS Sim 2 False 0.5 1.15 0.722
Qwen2-7B-Instruct FS Sim 4 False 0.5 1.15 0.714
Table 9: Model Performance for RiddleSense (Part 1). The column Quant indicates whether the model is quantized or not.
Model Method Num.Ex Quant Temp Rep_Pen Average
RISCORE Results
Llama3-70B Fewshot for QA & Llama3-70B for distractors
Meta-Llama-3-70B-Instruct RISCORE Sim 2 True 0.5 1.15 0.792
Meta-Llama-3-70B-Instruct RISCORE Sim 8 True 0.5 1.15 0.789
Meta-Llama-3-70B-Instruct RISCORE Sim 4 True 0.5 1.15 0.783
Mixtral-8x7B-Instruct-v0.1 RISCORE Sim 8 False 0.5 1.0 0.700
Mixtral-8x7B-Instruct-v0.1 RISCORE Sim 4 False 0.5 1.0 0.689
Mixtral-8x7B-Instruct-v0.1 RISCORE Sim 2 False 0.5 1.0 0.672
Meta-Llama-3-8B-Instruct RISCORE Sim 4 False 0.5 1.0 0.722
Meta-Llama-3-8B-Instruct RISCORE Sim 8 False 0.5 1.0 0.708
Meta-Llama-3-8B-Instruct RISCORE Sim 2 False 0.5 1.0 0.692
Mistral-7B-Instruct-v0.2 RISCORE Sim 2 False 0.5 1.0 0.600
Mistral-7B-Instruct-v0.2 RISCORE Sim 4 False 0.5 1.0 0.600
Mistral-7B-Instruct-v0.2 RISCORE Sim 8 False 0.5 1.0 0.597
Qwen2-7B-Instruct RISCORE Sim 8 False 0.5 1.0 0.731
Qwen2-7B-Instruct RISCORE Sim 4 False 0.5 1.0 0.717
Qwen2-7B-Instruct RISCORE Sim 2 False 0.5 1.0 0.697
Llama3-70B Fewshot for QA & Llama3-8B for distractors
Meta-Llama-3-70B-Instruct RISCORE Sim 4 True 0.5 1.15 0.789
Meta-Llama-3-70B-Instruct RISCORE Sim 2 True 0.5 1.15 0.786
Meta-Llama-3-70B-Instruct RISCORE Sim 8 True 0.5 1.15 0.775
Mixtral-8x7B-Instruct-v0.1 RISCORE Sim 2 False 0.5 1.0 0.719
Mixtral-8x7B-Instruct-v0.1 RISCORE Sim 8 False 0.5 1.0 0.689
Mixtral-8x7B-Instruct-v0.1 RISCORE Sim 4 False 0.5 1.0 0.686
Meta-Llama-3-8B-Instruct RISCORE Sim 8 False 0.5 1.0 0.706
Meta-Llama-3-8B-Instruct RISCORE Sim 4 False 0.5 1.0 0.686
Meta-Llama-3-8B-Instruct RISCORE Sim 2 False 0.5 1.0 0.681
Mistral-7B-Instruct-v0.2 RISCORE Sim 8 False 0.5 1.0 0.617
Mistral-7B-Instruct-v0.2 RISCORE Sim 4 False 0.5 1.0 0.606
Mistral-7B-Instruct-v0.2 RISCORE Sim 2 False 0.5 1.0 0.603
Qwen2-7B-Instruct RISCORE Sim 8 False 0.5 1.0 0.719
Qwen2-7B-Instruct RISCORE Sim 4 False 0.5 1.0 0.697
Qwen2-7B-Instruct RISCORE Sim 2 False 0.5 1.0 0.681
Llama3-8B Fewshot for QA & Llama3-8B for distractors
Meta-Llama-3-70B-Instruct RISCORE Sim 8 True 0.5 1.15 0.806
Meta-Llama-3-70B-Instruct RISCORE Sim 2 True 0.5 1.15 0.792
Meta-Llama-3-70B-Instruct RISCORE Sim 4 True 0.5 1.15 0.778
Mixtral-8x7B-Instruct-v0.1 RISCORE Sim 4 False 0.5 1.0 0.714
Mixtral-8x7B-Instruct-v0.1 RISCORE Sim 8 False 0.5 1.0 0.689
Mixtral-8x7B-Instruct-v0.1 RISCORE Sim 2 False 0.5 1.0 0.681
Meta-Llama-3-8B-Instruct RISCORE Sim 4 False 0.5 1.0 0.700
Meta-Llama-3-8B-Instruct RISCORE Sim 2 False 0.5 1.0 0.689
Meta-Llama-3-8B-Instruct RISCORE Sim 8 False 0.5 1.0 0.686
Mistral-7B-Instruct-v0.2 RISCORE Sim 8 False 0.5 1.0 0.614
Mistral-7B-Instruct-v0.2 RISCORE Sim 4 False 0.5 1.0 0.600
Mistral-7B-Instruct-v0.2 RISCORE Sim 2 False 0.5 1.0 0.589
Qwen2-7B-Instruct RISCORE Sim 2 False 0.5 1.0 0.694
Qwen2-7B-Instruct RISCORE Sim 8 False 0.5 1.0 0.689
Qwen2-7B-Instruct RISCORE Sim 4 False 0.5 1.0 0.683
Table 10: Model Performance for RiddleSense (Part 2). The column Quant indicates whether the model is quantized or not.